The following equation summarizes the trend portion of quarterly sales of condominiums over a long cycle. Sales also exhibit seasonal variations.

Ft = 65 − 5.5t + 4.5t 2
  

where
Ft = Unit sales
t = 0 at the first quarter of last year

Quarter Relative
1 .95
2 .90
3 .45
4 1.70


Using the information given, prepare a forecast of sales for each quarter of next year (not this year), and the first quarter of the year following that. (Round intermediate calculations and final answers to 2 decimal places.)

Quarter Forecast
1
2
3
4
1
Answer #1

For next year, below is the data for t

Quarter 1, t = 8

Quarter 2, t = 9

Quarter 3, t = 10

Quarter 4, t = 11

Quarter 1 of following year, t = 12

Estimation for t = 8, F = 65-5.5*8+4.5*(8)^2 = 309

Estimation for t = 9, F = 65-5.5*9+4.5*(9)^2 = 380

Estimation for t = 10, F = 65-5.5*10+4.5*(10)^2 = 460

Estimation for t = 11, F = 65-5.5*11+4.5*(11)^2 = 549

Estimation for t = 12, F = 65-5.5*12+4.5*(12)^2 = 647

Now,

Forecast for t = 8, F = 309*.95 = 293.55

Forecast for t = 9, F = 380*.90 = 342

Forecast for t = 10, F = 460*.45 = 207

Forecast for t = 11, F = 549*1.70 = 933.30

Forecast for t = 12, F = 647*.95 = 614.65